oregonduck (oregonduck):

MEDAL AND FAN FOR ANSWER What is (3x+2)(x+1)

oregonduck (oregonduck):

it has to have three terms

zarkam21 (zarkam21):

What is the question?

oregonduck (oregonduck):

What is (3x+2)(x+1)

oregonduck (oregonduck):

@aloud

oregonduck (oregonduck):

@Greg_D

OpenStudy (theanarchist):

I got this gimme a sec first

oregonduck (oregonduck):

ok

OpenStudy (theanarchist):

3xSquared+5x+3

OpenStudy (anonymous):

(3x+2)(x+1) =(3x+2)(x+1) =(3x)(x)+(3x)(1)+(2)(x)+(2)(1) =3x2+3x+2x+2 =3x2+5x+2

OpenStudy (theanarchist):

HA I got it first

oregonduck (oregonduck):

THXS